OpenStudy (anonymous):

how do i solve this? 0.5n+17+n=20 step by step work

OpenStudy (anonymous):

Get n by itself. You need to subtract .5, and 17 from both sides of the equatioon

OpenStudy (anonymous):

Wait sorry disregard my post. I didnt see .5n

OpenStudy (anonymous):

Oh snap. Me too. still get rid of 17 and then divide by 1.5

OpenStudy (anonymous):

20-17=n+.5 3=1.5n N is = to 1 in this case n=2

OpenStudy (anonymous):

wht?

OpenStudy (merchandize):

0.5n+17+n=20 0.5n+n=20-17 1.5n=3 n=3/1.5 n=2

OpenStudy (anonymous):

i need help with 0.5n+17+n=20 step by step